OpenStudy (anonymous):

Can someone help me with this rational expression? I'm going to draw it in this link. Please!

OpenStudy (anonymous):



OpenStudy (anonymous):

those are A's

OpenStudy (anonymous):

\[(4a-3)=\frac{a+13}{a+1}\]?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

\[(4a-3)(a+1)= a+13\] \[4a^2+a-3=a+13\] \[4a^2=16\] \[a^2=4\] \[a=2\ or\ a=-2\]